OpenStudy (asylum15):

log5(x+3) + log5(x-4) = 2log5(x-1) Done it 4 times, and keep getting -13. How? :(

OpenStudy (anonymous):

it can't be -13 because you cannot take the log of a negative number

OpenStudy (asylum15):

Yeah, I figured. I turned the left handside into log(AB) and on the right handside, I moved the 2 over to square the (x-1) the bases are the same so I can just equate, but the x squared(s) cancel! :(

OpenStudy (anonymous):

start with \[\log\left((x+3)(x-4)\right)=\log\left((x-1)^2\right)\]

OpenStudy (cwtan):

(x+3)(x-4)=(x-1)^2

OpenStudy (anonymous):

ok then lets see \[(x+3)(x-4)=(x-1)^2\] \[x^2-x-12=x^2-2x+1\] \[-x-12=-2x-1\] \[x-12=-1\] \[x=11\] is what i get

OpenStudy (asylum15):

Damn, looks like I made a sign error! Thank god for that! Thank you so much Satellite. :)

OpenStudy (anonymous):

i am going to make a guess my guess is that you wrote \[(x-1)^2=x^2-1\] which is wrong just a guess though

OpenStudy (anonymous):

yw

OpenStudy (kinggeorge):

I'd recommend checking the answer again. If \(x=11\) then it implies that \(14\cdot7=100\) but this is obviously false.

OpenStudy (kinggeorge):

I'm getting \[(x+3)(x-4)=(x-1)^2\]\[x^2-x-12=x^2-2x+1\]\[-x-12=-2x+1\]\[x-12=1\]\[x=13\]

OpenStudy (asylum15):

Yeah, kinggeorge, i got a negative 13, so it was a sign error, thank you for the response. Is it possible to ask you something in here?

OpenStudy (kinggeorge):

Sure.

OpenStudy (cwtan):

Lol I think it is typing error

OpenStudy (cwtan):

He mistaken the -2x+1 as -2x-1

OpenStudy (cwtan):

btw KingGeorge is right~

OpenStudy (asylum15):

I'm a biomedical engineer, doing technological maths. I can do differentiation, integration of the utmost difficulty without hesitation but its these simpler equations I am trying to remember how to do (its been a while) There was a question I was given, that I was wondering if you could help with: Q. Suppose the area of a rectangle is 114.4m2 and the length of the rectangle is 14m longer that its width. Calculate the length and width of the rectangle. If i get that, I'm ready for my exam :)

OpenStudy (asylum15):

cwtan, an error from me?

OpenStudy (cwtan):

@satellite73 's typing error

OpenStudy (kinggeorge):

To find the area of that, you get an expression \[x(x+14)=114.4\]Where \(x\) is the width, and \(x+14\) is the length.

OpenStudy (asylum15):

Yeah, I did it right, and that becomes a quadratic? The 2 x values you get, you discard one and refill in for X in Width + Length?

OpenStudy (kinggeorge):

You would discard the negative x-value, and the other x-value would be your width.

OpenStudy (asylum15):

Fantastic! You've been a great help my friend. I guess its time I head over to the biology side of this site and help some people? :)

OpenStudy (kinggeorge):

You're welcome.